OpenStudy (anonymous):

Can someone please help me with these few math problems. 1) 6x-9x-4=2x-2 2) -(x-7)=6x+8 3) 3-6a=9-5a 4) -9+6=-x+4 I have alot of trouble with math plaese explain thank you!

OpenStudy (anonymous):

1) \[6x-9x-4=2x-2\] \[-3x-4=2x-2\] \[-5x=2\] \[\frac{ -5x }{ -5} = \frac{ 2 }{ -5 }\] \[x=-\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

-(x-7)=6x+8 multiply out the brackets so it becomes: -x+7=6x+8 then get all the x to one side and the numbers to the other so then it becomes: -x-6x=8-7 which then becomes: -7x=1 then devide both sides by -7 so your answer is positive which gives you: \[\frac{ -7x }{ -7}=\frac{ 1 }{ -7 }\] and which finally equals: \[x=-\frac{ 1}{ 7 }\]

OpenStudy (anonymous):

that was 2), can you do the next two?